# Fusing Bluetooth with Pedestrian Dead Reckoning: A Floor Plan-Assisted Positioning Approach

### v2

@@ V-B BLE Positioning Results
We first analyze the experimental results of the proposed GML algorithm. In terms of baselines, in addition to the traditional maximum likelihood (ML) algorithm in our previous work[1], we also conduct another two common algorithms: the linear least squares (LLS)-based trilateration[23] and the filtered RSSI and beacon weight (FRBW) algorithm[47]. Note that: (i) The LLS-based trilateration takes all beacons into calculation. (ii) For the FRBW algorithm, the weight adjustable degree $g$ is set to $5$, and only the top $3$ beacons with the largest RSSIs are involved in the calculation. (iii) We apply the same mean filtering used in GML algorithm to the results of all algorithms. Taking the step moments detected by FP-BP as the reference, we record the BLE estimated position at each step moment, and calculate its PE relative to the ground truth position.
The experimental results are shown in Fig. 11. It can be observed that despite using KF and mean filtering, the positioning results estimated by the trilateration still exhibit significant fluctuations, with some results deviating considerably from the ground truth. Both GML and FRBW algorithms successfully constrained their results within the experimental area. However, FRBW algorithm still has some estimated positions appearing inside obstacles, which is avoided by the proposed GML algorithm.
Figs. 11 and 11 present the comparison of performance metrics between GML and baselines. It can be seen that GML algorithm achieves an MPE of $2.01~{}\mathrm{m}$, which is the lowest among all baselines, while the MPEs of trilateration, FRBW, and ML algorithms are $9.39~{}\mathrm{m}$, $2.55~{}\mathrm{m}$, and $2.63~{}\mathrm{m}$, respectively. Moreover, in terms of $80\%$ CDF corresponding PE, the GML algorithm results in $2.90~{}\mathrm{m}$, whereas FRBW and ML result in $3.51~{}\mathrm{m}$ and $3.55~{}\mathrm{m}$, respectively. These demonstrate that the GML algorithm significantly outperforms these baselines in terms of positioning performance.
@@ V-C FP-BP Positioning Results
To assess the performance of FP-BP, we conduct the fusion positioning systems proposed by Jin et al.[48], Xia et al.[33], Wang et al.[37] and Xu et al.[39] as baselines for comparison. In particular: (i) The system proposed by Jin et al.[48] is a common BLE/PDR fusion system without floor plan integration. (ii) Xia et al.[33] introduce the floor plan into the traditional PF-based BLE/PDR fusion system. In their system, the weights of particles within obstacle areas are reset to zero. (iii) Wang et al.[37] identify landmarks such as access points, doors, corners, and walls in the floor plan, and implement specify position calibrations in their system. Note that due to the complexity of this floor plan, it is challenging to define the direction of corners. Therefore, only the positions of the corners are matched in our experiments. (iv) Xu et al.[39] proposed an optimization scheme for particle re-initialization based on[33], taking into account the problem of particle depletion and result divergence. Note that we utilize the KF-processed BLE RSSIs in place of the WiFi fine timing measurement (FTM) due to the challenges in implementing time-based BLE positioning[13].
@@
The experimental results are shown in Figs. $1$, and 12. Fig. 11 demonstrates the experimental results of FP-BP in different cases, including FP-BP, FP-BP without the PPC algorithm (denoted as BP), and PDR. It is evident that although the initial position of PDR is set as the ground truth, the PE of PDR gradually increases as sensor errors accumulate, eventually deviating significantly from the ground truth. After fusing the BLE estimated position using PF, the cumulative error of BP is significantly reduced; however, many positioning results still “pass through walls” and appear outside the intended walkable area. Finally, by applying the PPC algorithm to incorporate the floor plan, the FP-BP algorithm successfully confines these positioning points within the correct area.
Figs. 11 and 11 show the comparison of PE for FP-BP in different cases, including the proposed GML algorithm. In these cases, we can see FP-BP achieves the lowest MPE of $1.19~{}\mathrm{m}$, followed by BP with an MPE of $1.51~{}\mathrm{m}$. The system that uses PDR solely has the largest MPE of $7.41~{}\mathrm{m}$, indicating that the cumulative errors of IMU significantly increase as the user walks further. However, after incorporating BLE, the MPE drops noticeably, reaching below $2~{}\mathrm{m}$, demonstrating that BLE positioning effectively helps mitigate the cumulative error. Additionally, after integrating PDR, BP outperforms GML in terms of accuracy, suggesting that PDR provides valuable movement information that enhances BLE positioning. Finally, we observe that after applying the PPC algorithm, the MPE of FP-BP is further reduced by approximately $0.3~{}\mathrm{m}$, which clearly demonstrates that floor plan integration significantly improves the positioning accuracy by $20\%$. Additionally, for FP-BP, the 50% and 80% CDF corresponding PEs are $1.13~{}\mathrm{m}$ and $1.60~{}\mathrm{m}$, respectively, representing improvements of $17\%$ and $26\%$ compared to BP. Overall, FP-BP achieves the best performance among all these cases.
@@
To further evaluate the performance of FP-BP, in Figs. $1$ and 12, we take the baseline systems[37, 33, 39],[48] for comparison. In Fig. 11, we can see that due to the lack of floor plan integration, some estimated positions of[48] appear inside obstacles. Additionally, despite the incorporation of the floor plan, baseline[39] still cannot prevent some results from erroneously appearing within obstacles. Other algorithms, including FP-BP, have successfully avoided this “wall-crossing” problem.
Fig. 11 shows the CDFs of PE for FP-BP and baselines. We can observe that among these algorithms, FP-BP achieves the lowest MPE at $1.19~{}\mathrm{m}$, consistent with Fig. 11. In contrast, the MPEs of the baselines are $2.20~{}\mathrm{m}$[48], $1.67~{}\mathrm{m}$[33], $2.02~{}\mathrm{m}$[37], and $1.97~{}\mathrm{m}$[39], respectively, which are significantly larger that of FP-BP, and FP-BP improves positioning accuracy by over $28\%$ compared with these baselines. In Fig. 11, we can further observe that in terms of 80% CDF corresponding PE, the FP-BP algorithm results in $1.60~{}\mathrm{m}$, while the PEs of these baselines are all over $2~{}\mathrm{m}$.
Fig. 12 demonstrates the PEs for FP-BP and baseline algorithms as the step count increases in a single experiment. It is evident that all these algorithms effectively reduce cumulative errors, as their PEs do not significantly increase with the step count. However, the results of[48, 33, 37, 39] exhibit significant fluctuations, with peak PEs exceeding $3~{}\mathrm{m}$, indicating that these algorithms struggle to provide accurate and stable positioning in certain areas. In contrast, the PE of FP-BP remains more stable, with about $90\%$ of PEs kept within $2~{}\mathrm{m}$, and no instances exceeding $3~{}\mathrm{m}$. These results indicate that FP-BP demonstrates superior performance compared to existing fusion-based positioning systems.
@@ V-D Results Comparison Across Trajectories
TABLE II: Experimental Results of FP-BP
Trajectory MPE [m] 50%CDF [m] 80%CDF [m] MCT [ms]
No. Steps FP-BP BP FP-BP BP FP-BP BP GML Fusion
1 528 1.19 1.51 1.13 1.32 1.60 2.14 5.57 4.67
2 229 1.08 1.25 1.01 1.22 1.48 1.56 4.88 5.35
3 313 0.89 1.18 0.80 1.07 1.33 1.58 5.23 4.58
@@
To verify the robustness of FP-BP under various trajectories, we further conduct experiments to compare the results of FP-BP across three trajectories, as shown in Figs. 13, 14 and Table II. In Fig. 13, we can see that similar to Fig. 11, FP-BP effectively constrains the positioning results within the correct Room. Moreover, when the user passes through a Door into another Room, it accurately handles this Room switch event.
@@
In Fig. 14, we can observe that across the three trajectories, as the step count increases, FP-BP similarly effectively mitigates cumulative errors, keeping almost all the PEs within $2~{}\mathrm{m}$. Additionally, Table II indicates that FP-BP demonstrates excellent positioning performance across these three trajectories, achieving MPEs of $1.08~{}\mathrm{m}$ and $0.89~{}\mathrm{m}$ for trajectories 2 and 3, respectively. In contrast, the BP system without floor plan integration reaches MPEs of $1.25~{}\mathrm{m}$ and $1.18~{}\mathrm{m}$ for the same trajectories. Furthermore, the 50% and 80% CDF corresponding PEs of FP-BP are also less than those of BP. These results indicate that FP-BP operates stably in indoor environments, remaining unaffected by changes in trajectory, and exhibits both high accuracy and strong robustness.
Finally, we analyze the computational delay of FP-BP to evaluate the system’s complexity. On the Android platform (see Table I for smartphone parameters), the GML and Fusion modules are implemented as independent Runnable tasks respectively, and each of them is executed at a fixed interval. Therefore, we separately report the MCT of these modules within FP-BP. From Table II, it can be seen that both these modules keep the MCT within $6~{}\mathrm{ms}$ for all three trajectories, which are far less than the BLE signal broadcasting interval and position estimation interval. This clearly demonstrates that the FP-BP system has strong real-time performance, low latency, and maintains good robustness through various experiments.
@@ VI Conclusion
In this paper, we have proposed the FP-BP algorithm, which can fully utilize the floor plan information to achieve real-time and accurate indoor positioning, without any limitations on indoor layout or map geometry. We have also proposed the GML algorithm to enhance RSSI-based BLE positioning. In particular, based on the accurate extraction of map features, a floor plan preprocessing method is first performed in the offline phase to construct an indoor map. In the online phase, based on an MLE model, we apply the GML algorithm to obtain the BLE result. Then, smartphone sensor data is fused by a PF to perform an initial estimation. Finally, based on the preprocessed indoor map, floor plan prior information is integrated through the proposed PPC algorithm by applying specific post-correction to the initial estimation. To verify the performance of FP-BP, we have conducted experiments in the public area with a total area of over 5,000 m2 in a building, and the results indicate that the FP-BP algorithm can achieve a mean positioning accuracy of 1.19 m, outperforming other floor plan-fused baselines by over 28%. Moreover, FP-BP algorithm has also demonstrated superior real-time performance on mobile devices. Therefore, it holds great promise for positioning in large-scale indoor scenarios.

### v3

@@ V-B BLE Positioning Results
Figure 12: Experimental results of BLE positioning under normal (dense) and sparse beacon deployments. “(S)” represents the sparse deployment.
We first analyze the experimental results of the proposed GML algorithm. In terms of baselines, in addition to the traditional maximum likelihood (ML) algorithm in our previous work[1], we also conduct another three common algorithms: a grid-based improved MLE (GIMLE) algorithm[41], the linear least squares (LLS)-based trilateration[23], and a filtered RSSI and beacon weight (FRBW) algorithm[49]. Note that: (i) The LLS-based trilateration takes all beacons into calculation. (ii) For the FRBW algorithm, the weight adjustable degree $g$ is set to $5$, and only the top $3$ beacons with the largest RSSIs are involved in the calculation. (iii) We apply the same mean filtering used in GML algorithm to the results of all algorithms. Taking the step moments detected by FP-BP as the reference, we record the BLE estimated position at each step moment, and calculate its error relative to the ground truth position.
The results are shown in Figs. 1111(a) and 11(b). Note that, since the accuracy of trilateration is significantly lower than that of the other methods, its trajectory is not presented in Fig. 1111(a). We can observe that despite using KF and mean filtering, the positioning results estimated by GIMLE still exhibit significant fluctuations, with some results deviating considerably from the ground truth. This is mainly because GIMLE does not take into account the geometric constraints formed by beacon positions. In contrast, both GML, ML and FRBW algorithms successfully constrained their results within the experimental area. However, FRBW algorithm still has some estimated positions appearing inside obstacles, which is avoided by the proposed GML algorithm.
Fig. 1111(b) and Table II present the comparison of performance metrics between GML and baselines. We can see that GML algorithm achieves an MPE of $1.74\,\mathrm{m}$, which is lower than all the baselines, with a significant error reduction of over $34\%$. In addition, GML significantly outperforms all the baselines in terms of P50/P80 scores, and STD. These results demonstrate that the proposed GML algorithm achieves higher positioning accuracy and better stability.
In addition to the normal (dense) beacon deployment, we also consider a sparse deployment scenario in Fig. 9. The performance under both deployments is compared in Fig. 12, using the CDFs of position error. Since ML and FRBW achieve the best performance in Table II, we select them as baselines. As shown, all algorithms experience degraded accuracy under the sparse deployment; however, GML still achieves the highest precision, with an MPE of $3.46\,\mathrm{m}$, outperforming ML ($5.76\,\mathrm{m}$) and FRBW ($4.77\,\mathrm{m}$). This demonstrates that GML maintains robust even in sparse layouts, since the constraint $\bm{y}\in\mathrm{int}\mathcal{C}_{N}$ in (22), which is unsuitable for sparse deployments, is removed to avoid large deviations.
TABLE II: Experimental Results
Methods MPE [m] P50 [m] P80 [m] STD [m]
GML (Ours) 1.74 1.53 2.61 1.08
GIMLE[41] 3.30 3.16 4.71 1.57
Trilateration[23] 9.45 4.54 11.89 15.48
FRBW[49] 2.64 2.21 3.90 1.99
ML[1] 2.68 2.17 3.55 2.05
FP-BP (Ours) 1.14 1.09 1.62 0.59
BP 1.46 1.34 2.08 0.91
PDR 7.47 6.81 12.07 5.02
Kong et al.[26] 2.32 2.08 3.36 1.36
Xia et al.[33] 1.61 1.41 2.25 1.11
Wang et al.[37] 2.02 1.74 2.72 1.22
Xu et al.[39] 1.97 1.42 2.48 1.91
@@ V-C FP-BP Positioning Results
To assess the performance of FP-BP, we conduct the fusion positioning systems proposed by Kong et al.[26], Xia et al.[33], Wang et al.[37] and Xu et al.[39] as baselines for comparison. In particular: (i) The system proposed by Kong et al.[26] is a tightly-coupled BLE/PDR fusion system without floor plan integration. (ii) Xia et al.[33] introduce the floor plan into the traditional PF-based BLE/PDR fusion system. In their system, the weights of particles within obstacle areas are reset to zero. (iii) Wang et al.[37] identify landmarks such as access points, doors, corners, and walls in the floor plan, and implement specify position calibrations in their system. Note that due to the complexity of this floor plan, it is challenging to define the direction of corners. Therefore, only the positions of the corners are matched in our experiments. (iv) Xu et al.[39] proposed an optimization scheme for particle re-initialization based on[33], taking into account the problem of particle depletion and result divergence. Note that we utilize the KF-processed BLE RSSIs in place of the WiFi fine timing measurement (FTM) to ensure a consistent comparison.
@@
The results are shown in Figs. 1111(c)11(f), 13, 14 and Table II. Figs. 1111(c) and 11(d) are the ablation experiments which compare the results of FP-BP, FP-BP without PPC (denoted as BP), and PDR. It is evident that although the initial position of PDR is set as the ground truth, the error of PDR gradually increases as sensor errors accumulate, eventually deviating significantly from the ground truth. After fusing the BLE estimated position using PF, the cumulative error of BP is significantly reduced; however, many positioning results still “pass through walls” and appear outside the intended walkable area. Finally, by applying the PPC algorithm to incorporate the floor plan, the FP-BP algorithm successfully confines these positioning points within the correct area.
In Fig. 1111(d) and Table II, we can observe that the original PDR yields the largest MPE of $7.47\,\mathrm{m}$. After fusing BLE results, the MPE of BP is significantly reduced to $1.46\,\mathrm{m}$. With the further introduction of the proposed PPC for post-processing, the MPE of FP-BP is further reduced to $1.14\,\mathrm{m}$, representing an improvement of about $22\%$. In addition, BP fused with PDR also achieves higher accuracy than the pure BLE-based GML algorithm. In terms of P50/P80 scores and STD, FP-BP demonstrates the best positioning accuracy and stability. These results validate the effectiveness of each component in the proposed FP-BP fusion system.
Figure 14: Comparison of MPE and MCT for different particle numbers.
To further evaluate the performance of FP-BP, in Figs. 1111(e), 11(f) and 13, we take the baseline systems[26, 37, 33] and[39] for comparison. In Fig. 1111(e), we can see that due to the lack of floor plan integration, some estimated positions of[26] appear inside obstacles. Additionally, despite the incorporation of the floor plan, baseline[39] still cannot prevent some results from erroneously appearing within obstacles. Other algorithms, including FP-BP, have successfully avoided this “wall-crossing” problem.
From Fig. 1111(f) and Table II, we can observe that among these algorithms, FP-BP achieves the lowest MPE at $1.14\,\mathrm{m}$, consistent with Fig. 1111(d). In contrast, the MPEs of other baselines are all over $1.5\,\mathrm{m}$, which are significantly larger that of FP-BP, and FP-BP improves positioning accuracy by over $29\%$ compared with these baselines. Moreover, compared with the other baselines, FP-BP improves P50, P80, and STD by at least $23\%$, $28\%$, and $47\%$, respectively, demonstrating superior positioning performance.
@@
Fig. 13 compares the errors for FP-BP and baseline algorithms as the step count increases in a single experiment. We can observe that all these algorithms effectively reduce cumulative errors, as their errors do not significantly increase with the step count. However, the results of[26, 33, 37] and[39] exhibit significant fluctuations, with peak errors exceeding $3\,\mathrm{m}$, indicating that they struggle to provide stable positioning in certain areas. In contrast, the errors of FP-BP remain more stable, with about $90\%$ of errors kept within $2\,\mathrm{m}$. These results indicate that FP-BP demonstrates superior robustness compared to other fusion-based systems.
In Fig. 14, we compare the MPE and MCT of FP-BP under different particle numbers. We can observe that as the particle number increases, the MPE decreases, which is consistent with (37), since more particles can enable a more accurate estimation of posterior PDF $p(\bm{x}|\mathcal{Z})$. Meanwhile, the MCT increases with the particle number. When the particle number $m>500$, the MPE converges to $1.14\,\mathrm{m}$; therefore, we set $m=500$ and the corresponding MCT $6\,\mathrm{ms}$. Furthermore, the proposed FP-BP requires significantly fewer particles than methods such as[38] and[39], since the PPC mechanism only corrects the particle positions rather than discarding them, thus avoiding particle loss.
@@ V-D Results Comparison Across Trajectories
TABLE III: Experimental Results Across Trajectories
Trajectory MPE [m] P50 [m] P80 [m]
No. Steps Dist. [m] FP-BP BP FP-BP BP FP-BP BP
1 528 316.8 1.14 1.46 1.09 1.34 1.62 2.08
2 229 137.4 0.93 0.98 0.92 0.95 1.26 1.38
3 313 187.8 0.80 1.13 0.74 0.98 1.23 1.49
4 193 115.8 1.07 2.13 1.12 1.65 1.49 3.62
@@
To verify the robustness of FP-BP under various trajectories, we further conduct comparative experiments across another three trajectories, as shown in Figs. 15, 16 and Table III. In Figs. 1515(a) and 15(b), we can see that similar to Fig. 1111(d), FP-BP effectively constrains the results within the correct Room. Moreover, when the user passes through a Door into another Room, it accurately handles this Room switch event. Fig. 1515(c) shows the results of cross-floor trajectory 4. We can see that the extended FP-BP can effectively handle multi-floor scenarios, and no additional error is introduced during floor transitions. This is because the algorithm is re-initialized upon reaching the new floor, thereby eliminating the accumulated error from the previous floor.
@@
In Fig. 16, we can observe that across all trajectories, as the step count increases, FP-BP similarly effectively mitigates cumulative errors, keeping almost all the errors within $2\,\mathrm{m}$. Additionally, Table III indicates that FP-BP demonstrates excellent positioning performance across trajectories 2–4, achieving MPEs of $0.93\,\mathrm{m}$, $0.80\,\mathrm{m}$ and $1.07\,\mathrm{m}$ respectively. In contrast, the BP system without floor plan integration reaches MPEs of $0.98\,\mathrm{m}$, $1.13\,\mathrm{m}$ and $2.13\,\mathrm{m}$. Furthermore, the P50 and P80 scores of FP-BP are also less than those of BP. These results indicate that FP-BP operates stably in indoor environments, remaining unaffected by changes in trajectory, and exhibits both high accuracy and strong robustness.
V-E Results Comparison Across Smartphones
TABLE IV: Smartphone Parameters
No. Model Mobile processor Operating system RAM
1 HONOR Magic 6 Snapdragon 8 Gen3 Android 15 12 GB
2 HUAWEI Mate 40 Pro Kirin 9000 Android 10 8 GB
3 realme Q2i Dimensity 720 Android 10 4 GB
Figure 17: Experimental results of FP-BP across smartphones.
TABLE V: Experimental Results Across Smartphones
No. MPE [m] P50 [m] P80 [m] STD [m] MCT1 [ms] MCT2 [ms]
1 1.14 1.09 1.62 0.59 5.57 4.67
2 1.57 1.36 2.38 0.95 2.76 6.31
3 1.63 1.50 2.27 0.82 5.94 7.67
The performance of PDR-based algorithms is typically closely related to the quality of smartphone sensors. To verify the feasibility of FP-BP across different smartphones, in addition to smartphone 1 in Table I, two additional smartphones shown in Table IV were introduced. These smartphones are equipped with different mobile processors and random access memories (RAMs). We conducted experiments along trajectory 1 using each of these smartphones, and the results are presented in Fig. 17 and Table V. The trajectories in Fig. 17 indicate that the proposed FP-BP algorithm yields consistent results across all three smartphones without significant deviation. The statistical results in Table V show that FP-BP performs best on smartphone 1. Although its performance slightly decreases on smartphones 2 and 3, the MPE remains around $1.5\,\mathrm{m}$, and the STD is below $1\,\mathrm{m}$, suggesting that the accuracy is stable and consistent across different positions.
In addition, in Table V, let MCT1 and MCT2 denote the MCT of GML module and PDR&Fusion module, respectively. We can observe that all MCTs are at the millisecond level, which are far below the BLE signal broadcasting and position estimation intervals. This demonstrates that the proposed FP-BP algorithm imposes a low computational burden on the device, offering strong feasibility for achieving low-latency real-time positioning.
@@ VI Conclusion
In this paper, we have proposed the FP-BP algorithm, which can fully utilize the floor plan information to achieve real-time fusion positioning without limitations on map geometry. We have also proposed the GML algorithm to enhance RSSI-based BLE positioning. In particular, based on the accurate extraction of map features, the floor plan is preprocessed in the offline phase to construct an indoor map. In the online phase, based on the MLE principle, we apply the GML algorithm to obtain BLE results. Then, smartphone sensor data is fused by a PF to perform an initial estimation. Finally, based on the preprocessed indoor map, floor plan information is integrated by applying the proposed PPC mechanism to the initial estimation. To evaluate the performance of FP-BP, we conducted both single- and cross-floor experiments in a building with a total area of over 5,000 m2 per floor. The results indicate that the FP-BP algorithm can achieve a mean positioning accuracy of 1.14 m, outperforming other floor plan-fused baselines by over 29%. Moreover, FP-BP algorithm has also demonstrated superior real-time performance on mobile devices. Therefore, it holds great promise for large-scale indoor positioning. In future work, we aim to extend FP-BP to support diverse user carrying poses and motion patterns, further improving its adaptability to real-world scenarios.